# Enantiomer-Selective Magnetoresistance in Chiral Gold Nanocrystals by Magnetic Control of Surface Potentials

### v1

@@ Abstract
Chiral nanomaterials offer intriguing possibilities for novel electronic and chemical applications. Here, we report the discovery of an enantiomer-selective magnetoresistance effect in chiral gold nanocrystals. Based on precise control of nanocrystal chiral morphology using amino acid-directed synthesis, we demonstrate that an external magnetic field can dramatically modulate resistance in an enantiomer-specific manner. For a given enantiomer, a magnetic field in one direction alters the resistance by over an order of magnitude, while the opposite field direction leaves it unchanged. This asymmetric response reverses for the opposite enantiomer. We attribute this phenomenon to a novel chirality-driven charge trapping mechanism, where the interplay between the chiral nanocrystal morphology and the magnetic field selectively modifies the surface potential. The magnitude and sign of the magnetoresistance can be further tuned by the surface chemistry of the nanocrystal, as demonstrated through sulfide treatment. Our findings reveal a new form of chirality-dependent magnetoresistance, distinct from previously known effects such as chirality-induced spin selectivity and electric magnetochiral anisotropy. The ability to remotely control surface potentials of chiral nanostructures using magnetic fields could enable novel approaches in catalysis, drug delivery, and nanoelectronics.
@@ 1 Introduction
In this work, we discovered a unique chirality-driven MR in chiral Au NCs by applying an external magnetic field. We synthesized Au NCs with pinwheel-like chiral morphology by an amino-acid-directed strategy and measured their resistance using conducting atomic force microscopy (cAFM). Distinct from the known CISS or EMCA, the resistance of chiral Au NCs is largely enhanced (positive MR) in one field orientation and nearly unchanged in the opposite field, showing a salient asymmetric feature with respect to the zero-field case. The asymmetric order was reversed when we switched to NCs with opposite handedness, verifying the role of chirality. By modifying the NC surface capping layer with sulfide, we can also reduce the resistance (negative MR) by one field direction and keep it unchanged in the opposite field. We rationalize the usual MR by the magnetochiral-driven charge trapping effect [13] in the surface capping layer, which controls the surface potential and the total resistance. Our work provides valuable insights to understand chirality-driven transport effects and offers implications for selectively controlling the surface potential and manipulate the surface activity for chemical and biological reactions.
@@ 2 Synthesis and characterization of chiral Au NCs
The Au NCs with intrinsic chirality were synthesized based on an amino-acid-directed chirality transfer strategy [23, 30] (see Methods and Figure 1). With L- or D-cysteine (Cys) as chiral inducers, spherical Au seeds were grown into chiral Au NCs with opposite chirality. After the syntheses, the Au NCs were collected by centrifugation and redispersed in water with cetyltrimethylammonium bromide (CTAB) as the stabilizing agent. Note that CTAB (achiral) is adsorbed on Au surfaces to form a core-shell structure [33]. The chiral Au NCs synthesized with L- and D-Cys exhibit inverted circular dichroism (CD) spectra with respect to each other, with dissymmetry factors (g-factors) of approximately -0.033 and 0.041, respectively (Figure 1b and Extended Data Fig. 1). From the scanning electron microscopy (SEM) images, the pyramid-like subunits on the chiral Au NCs are rotated clockwise or counterclockwise with respect to the achiral trisoctahedral Au NCs obtained with racemic Cys (Figures  1c,1f,1i,1l, and Extended Data Fig. 2). The rotation direction of the pyramidal subunits is dictated by the chirality of the chiral inducers, due to the enantioselective interactions of L- or D-Cys with R- or S- chiral Au surfaces [23, 30]. After the rotation, the four adjacent pyramids on the trisoctahedron no longer share one vertex, but instead form a pinwheel-like structure. As shown in Figures  1d and 1m, the pinwheel-like structure of the chiral Au NCs is rotated clockwise and counterclockwise, respectively. To clarify, the Au NCs obtained with L- and D-Cys are designated as L-Au and D-Au NCs, respectively, and the achiral Au NCs with racemic Cys is designated as A-Au NCs (Extended Data Table 1). Furthermore, by adjusting the amounts of seeds in the growth processes, chiral Au NCs can be synthesized with sizes ranging from 150.2 to 475.5 nm with tunable chiroplasmonic properties (Extended Data Figs. 3,  4, and Extended Data Table 2).
@@ 3 Magnetoresistance of chiral Au NCs
Figure 2: Magneto-transport and asymmetric MR for chiral Au NCs. (a) An AFM image of D-Au NCs on the indium tin oxide (ITO) substrate. The scale bar is 1 $\mu$m. (b) Schematic of the ITO/chiral Au NC/Pt junction constructed via cAFM. Take the D-Au NC as an example. A permanent magnet (300 Oe) is placed beneath the ITO substrate. (c) and (d) The experimental (dots) and fitted (line, according to Eq. (1)) current-voltage curves for the L-Au NC and D-Au NC under different directions of magnetic fields, where the south pole of the magnet is pointing perpendicularly toward (+B, red) and away from (–B, blue) the ITO plane. 0B (black) stands for zero magnetic field. (e) and (f) The MR of the L-Au NC and D-Au NC under different directions of an external magnetic field.
@@
We measured the MR of the chiral Au NCs by cAFM with a Pt tip at room temperature. Using chiral Au NCs with size of $\sim$150 nm as exemplars, we dispersed L- and D-Au NCs on an indium tin oxide (ITO) substrate, enabling MR measurements at the single-nanoparticle level (Figure 2). The size of the L- and D-Au NC is $150.2\pm 4.8$ and $154.4\pm 3.6$ nm, respectively. We recorded current-voltage (I-V) relationships of individual chiral Au NCs via the ITO/chiral Au NC/Pt junction. An external magnetic field (B = 300 Oe) was applied orthogonally to the ITO substrate using a permanent magnet placed below the ITO (Figure 2b). Unlike CISS MR studies [34], our setup did not incorporate a magnetic substrate or tip. The insulating CTAB molecular shell on Au NCs contributes significantly more to the resistance than the metallic Au NC core [35], resulting in small currents of tens of nA at 1.0 V bias. Figures 2c and 2d present the averaged I-V curves for L- and D-Au NCs, respectively, derived from hundreds of individual measurements (see Extended Data Fig. 5). The I-V curves were measured 320–679 times on more than 30 individual Au NCs from at least three parallel samples. The error bar was calculated as the standard deviation of all results. We observed an asymmetric MR behavior between +B and –B (Figure 2). For L-Au NCs, the I-V curve at +B closely resembles the zero-field case within error bars, while at –B, the resistance increases dramatically by 20–50 times. Conversely, for D-Au NCs, –B barely affects the resistance, while +B increases it by $\sim$20 times. This reversal of asymmetric behavior between L- and D-Au NCs strongly indicates its chirality-dependent origin. We quantify MR as the relative change in resistance due to the B-field: MR=$[R(\textbf{B})-R(0)]/R(0)$. The results show that MR is negligible at +B (–B) but substantial at –B (+B) for L(D)-Au NCs, respectively (Figures 2e and 2f).
@@
To determine whether the chiral morphology of the Au core or the chiral ligand shell is primarily responsible for the observed MR results, we conducted control experiments using three additional types of Au NCs. In addition to the L-Au and D-Au NCs, we investigated achiral Au NCs synthesized with racemic Cys (A-Au, Extended Data Fig.2), L-cysteine-modified achiral Au NCs (Lm-A-Au), and D-cysteine-modified achiral Au NCs (Dm-A-Au). We prepared Lm-A-Au and Dm-A-Au by modifying achiral non-Cys Au NCs with L- and D-Cys, respectively (Extended Data Figs.6-7). Our initial measurements of the I-V behavior of A-Au NCs revealed negligible MR for ±B (Figures 3a, 3d, and Extended Data Fig. 8). For Lm-A-Au and Dm-A-Au NCs, while the asymmetric order between ±B showed a trend similar to L- and D-Au NCs, the magnitude of their MR was significantly smaller (Figure 3). The I-V curves under ±B and zero-field conditions were only marginally or hardly distinguishable when considering error bars. These findings strongly suggest that the substantial MR observed in chiral Au NCs is primarily attributable to the chiral morphology of the NC core rather than the molecular chirality of the ligand shell. This conclusion highlights the crucial role of the NC’s intrinsic morphology chirality in determining its MR properties.
@@
Figure 3: Magnetotransport of A-Au, Lm-A-Au, and Dm-A-Au NCs. (a)-(c) Current-voltage curves of the A-Au, Lm-A-Au, and Dm-A-Au NCs (from left to right) under magnetic field directions of +B (red), –B (blue), and 0B (black), respectively. At least 430 current-voltage curves were obtained for all samples, with approximately 10 measurements performed on each Au NC. The error bar was defined as the standard deviation of all current values measured at each magnetic field condition.
@@
Having established the chiral origin of the observed MR, we proceeded to investigate the robustness of this chirality-driven effect under various conditions, including different NC sizes, magnetic field strengths, and AFM forces, to rule out potential artifacts. Using L-Au NCs as a model system, we observed that the current through chiral Au NCs increases gradually as the NC size increases from 150.2 to 350.5 nm, consistent with previously reported transport behaviors of metal nanoparticles [35] (Extended Data Figs. 9a and 10). Concurrently, the overall MR value decreases with increasing NC size, while the MR values under –B consistently remain significantly larger than those under +B (Extended Data Fig. 9d), confirming the persistence of asymmetric MR across different NC sizes. The magnitude of the asymmetric MR also varies proportionally with the applied magnetic field strength. For L-Au NCs, as the external magnetic field decreases from 300 to 3 Oe, the current remains nearly constant for +B, while it increases from 0.5 to 21 nA for –B at a bias of -1.0 V. This corresponds to a substantial MR drop from 5100% to 18% (Extended Data Figs. 9b, 9e, and 11). In cAFM measurements, the contact between the AFM tip and Au NCs significantly influences the total resistance [36, 37]. To exclude potential artifacts arising from this contact, we evaluated the MR of L-Au NCs under various AFM forces. As anticipated, the current values increase with incremental force due to the compression of the molecular layer and the increased contact area between the tip and the Au NC. Notably, the asymmetric MR remains robust across different applied forces, although the MR at –B exhibits a weakly nonmonotonic change with increasing force (Extended Data Figs. 9c and 9f). This robustness against varying AFM force may be attributed to the stable morphology of crystalline nanoparticles compared to organic materials. These comprehensive investigations across different experimental parameters underscore the intrinsic nature and reliability of the observed chirality-driven magnetoresistance effect in Au NCs.
@@ 4 Control the sign of magnetoresistance
Figure 4: Magnetoresistance characterization of the sulfide-modified chiral Au NCs. (a) and (b) The experimental (dots) and fitted (line) current-voltage curves of Sm-L-Au and Sm-D-Au NCs at the magnetic field directions of +B (red), –B (blue), and 0B (black). We measured 180–650 current-voltage curves from more than 20 individual Au NCs and estimated the error bar by the standard deviation. (c) and (d) The MR of Sm-L-Au and Sm-D-Au NCs changing with bias under different external magnetic field directions. The Sm-L-Au and Sm-D-Au NCs were prepared by treating the L-Au and D-Au NCs with sodium sulfide, respectively.
@@
We discovered that the sign of asymmetric MR is sensitive to the chemical composition of the surface molecular layer on chiral Au NCs. We further treated the chiral Au NCs with sodium sulfide, resulting in additional sulfide adsorption into the CTAB layers on the Au NCs [38]. For clarity, we designate the sulfide-modified L- and D-Au NCs as Sm-L-Au and Sm-D-Au NCs, respectively. X-ray photoelectron spectroscopy studies confirm the presence of sulfide on their surfaces (Extended Data Fig. 12). SEM and spectra studies demonstrate that Sm-L-Au and Sm-D-Au NCs maintain similar chiral morphologies and optically chiral properties to their non-modified counterparts (Extended Data Figs. 13 and 14). Dynamic light scattering reveals slight size decreases for Sm-L-Au and Sm-D-Au NCs, attributed to the partial desorption of CTAB upon sulfide modification (Extended Data Fig. 15 and Extended Data Table 3).
@@
We investigated the MR of Sm-L-Au and Sm-D-Au NCs using the same cAFM setup (Figure 4 and Extended Data Fig. 16). At zero magnetic field, sulfide-modified NCs exhibit significantly higher conductance than pristine NCs, likely due to the thinner CTAB molecular layer. Despite this conductance change, the asymmetric nature of MR persists. For Sm-L(D)-Au NCs, –B (+B) field dramatically reduces the resistance by several times, while +B (–B) barely affects it. In contrast to L(D)-Au NCs, Sm-L(D) NCs display negative MR, as shown in Figure 4. This sign change in MR suggests that the magnetotransport properties are intimately related to the characteristics of the molecular shell layer.

### v2

@@ Abstract
Chiral nanomaterials offer intriguing possibilities for novel electronic and chemical applications. Here, we report the discovery of an enantiomer-selective magnetoresistance effect in chiral gold nanocrystals. Based on precise control of nanocrystal chiral morphology using amino acid-directed synthesis, we demonstrate that an external magnetic field can dramatically modulate resistance in an enantiomer-specific manner. For a given enantiomer, a magnetic field in one direction alters the resistance by dozens of times, while the opposite field direction leaves it unchanged. This asymmetric response reverses for the opposite enantiomer and are reproduced in both single nanocrystals by conduction atomic force microscopy and nanocrystal thin films in solid state devices. We attribute this phenomenon to a chirality-driven charge pumping effect, where the interplay between the chiral morphology and the magnetic field selectively modifies the surface potential. The magnitude and sign of the magnetoresistance can be further tuned by the surface chemistry of the nanocrystal, as demonstrated through sulfide treatment. Our findings reveal a new form of chirality-dependent magnetoresistance, distinct from previously known effects such as chirality-induced spin selectivity and electric magnetochiral anisotropy. The ability to remotely control surface potentials of chiral nanostructures using magnetic fields could enable novel approaches in catalysis, drug delivery, and nanoelectronics.
@@ 1 Introduction
In this work, we discovered a unique chirality-driven MR in chiral Au NCs by applying an external magnetic field. We synthesized Au NCs with pinwheel-like chiral morphology by an amino-acid-directed strategy and measured the resistance of single NCs using conducting atomic force microscopy (cAFM) and NC films by solid-state devices. Distinct from the known CISS or EMCA, the resistance of chiral Au NCs is largely enhanced (positive MR) in one field orientation and nearly unchanged in the opposite field, showing a salient asymmetric feature with respect to the zero-field case. The asymmetric order was reversed when we switched to NCs with opposite handedness, verifying the role of chirality and remains robust against temperature (10-400 K). By modifying the NC surface capping layer with sulfide, we can also reduce the resistance (negative MR) by one field direction and keep it unchanged in the opposite field. We rationalize the usual MR by the magnetochiral charge pumping [13] in the surface capping layer, which controls the surface potential and the total resistance. Our work provides valuable insights to understand chirality-driven transport effects and offers implications for selectively controlling the surface potential and manipulate the surface activity for chemical and biological reactions.
@@ 2 Synthesis and characterization of chiral Au NCs
The Au NCs with intrinsic chirality were synthesized based on an amino-acid-directed chirality transfer strategy [23, 30] (see Methods and Figure 1). With L- or D-cysteine (Cys) as chiral inducers, spherical Au seeds were grown into chiral Au NCs with opposite chirality. After the syntheses, the Au NCs were collected by centrifugation and redispersed in water with cetyltrimethylammonium bromide (CTAB) as the stabilizing agent. Note that CTAB (achiral) is adsorbed on Au surfaces to form a core-shell structure [33]. To clarify, the Au NCs obtained with L- and D-Cys are designated as L-Au and D-Au NCs, respectively, and the achiral Au NC with racemic Cys is designated as A-Au NCs (Extended Data Table 1) The chiral Au NCs synthesized with L- and D-Cys exhibit inverted circular dichroism (CD) spectra with respect to each other, with dissymmetry factors (g-factors) of approximately -0.033 and 0.041, respectively (Figure 1b and Extended Data Figure 1). From the scanning electron microscopy (SEM) images, the pyramid-like subunits on the chiral Au NCs are rotated clockwise or counterclockwise with respect to the achiral trisoctahedral Au NCs obtained with racemic Cys (Figures 1c,1f,1i,1l, and Extended Data Figure 2). The rotation direction of the pyramidal subunits is dictated by the chirality of the chiral inducers, due to the enantioselective interactions of L- or D-Cys with R- or S- chiral Au surfaces [23, 30]. After the rotation, the four adjacent pyramids on the trisoctahedron no longer share one vertex, but instead form a pinwheel-like structure. As shown in Figures 1d and 1m, the pinwheel-like structure of the chiral Au NCs is rotated clockwise and counterclockwise, respectively. Furthermore, by adjusting the amounts of seeds in the growth processes, chiral Au NCs can be synthesized with sizes ranging from 150.2 to 475.5 nm with tunable chiroplasmonic properties (Extended Data Figures 3 and  4, and Extended Data Table 2). We further characterized the magnetic properties of chiral and achiral Au NCs by superconducting quantum interference device (SQUID) measurements. Independent from chirality, these Au NCs exhibit extremely weak ferromagnetism (Extended Data Figure 5), similar to earlier observation in Au nanoparticles[34].
@@ 3 Magnetoresistance of chiral Au NCs
[…]
@@
We measured the MR of the chiral Au NCs by cAFM with a Pt tip at room temperature. Using chiral Au NCs with size of $\sim$150 nm as exemplars, we dispersed L- and D-Au NCs on an indium tin oxide (ITO) substrate, enabling MR measurements at the single-nanoparticle level (Figure 2). The size of the L- and D-Au NCs are $150.2\pm 4.8$ and $154.4\pm 3.6$ nm, respectively. We recorded current-voltage (I-V) relationships of individual chiral Au NCs via the ITO/chiral Au NC/Pt junction. An external magnetic field (B = 300 Oe) was applied orthogonally to the ITO substrate using a permanent magnet placed below the ITO (Figure 2b). Unlike CISS MR studies [35], our setup did not incorporate a magnetic substrate or tip. The insulating CTAB molecular shell on Au NCs contributes significantly more to the resistance than the metallic Au NC core [36], resulting in small currents of tens of nA at 1.0 V bias. Figures 2c and 2d present the averaged I-V curves for L- and D-Au NCs, respectively, derived from hundreds of individual measurements (see Extended Data Figure 6). The I-V curves were measured 320–679 times on more than 30 individual Au NCs from at least three parallel samples. The error bar was calculated as the standard deviation of all results. We observed an asymmetric MR behavior between +B and –B (Figure 2). For L-Au NC, the I-V curve at +B closely resembles the zero-field case within error bars, while at –B, the resistance increases dramatically by 20–50 times. Conversely, for D-Au NCs, –B barely affects the resistance, while +B increases it by $\sim$20 times. The I-V curves of L-Au NC display consistent behavior with those of ITO/L-Au NC/Pt when the Pt AFM tip is replaced with Au tip, indicating that the electrode material does not affect the MR behavior (Extended Data Figure 7). The I-V curves obtained through direct contact of ITO/Pt electrodes are independent of the magnetic field (Extended Data Figure 8). This reversal of asymmetric behavior between L- and D-Au NCs strongly indicates its chirality-dependent origin. We quantify MR as the relative change in resistance due to the B-field: MR=$[R(\textbf{B})-R(0)]/R(0)$. The results show that MR is negligible at +B (–B) but substantial at –B (+B) for L(D)-Au NCs, respectively (Figures 2e and 2f).
@@
To determine whether the chiral morphology of the Au core or the chiral ligand shell is primarily responsible for the observed MR results, we conducted control experiments using three additional types of Au NCs. In addition to the L-Au and D-Au NCs, we investigated achiral Au NCs synthesized with racemic Cys (A-Au, Extended Data Figure 2), L-cysteine-modified achiral Au NCs (Lm-A-Au), and D-cysteine-modified achiral Au NCs (Dm-A-Au). We prepared Lm-A-Au and Dm-A-Au by modifying achiral non-Cys Au NCs with L- and D-Cys, respectively (Extended Data Figures 9 and 10). Our initial measurements of the I-V behavior of A-Au NCs revealed negligible MR for ±B (Figures 3a and Extended Data Figure 11). For Lm-A-Au and Dm-A-Au NCs, while the asymmetric order between ±B showed a trend similar to L- and D-Au NCs, the magnitude of their MR was significantly smaller (Figure 3). The I-V curves under ±B and 0B were only marginally or hardly distinguishable when considering error bars for A-, Lm- and Dm-Au NCs. These findings strongly suggest that the substantial MR observed in chiral Au NCs is primarily attributable to the chiral morphology of the NC core rather than the molecular chirality of the ligand shell. This conclusion highlights the crucial role of the NC’s intrinsic morphology chirality in determining its MR properties.
@@
Figure 3: Magnetotransport of A-Au, Lm-A-Au, and Dm-A-Au NCs. (a)-(c) Current-voltage curves of the A-Au, Lm-A-Au, and Dm-A-Au NCs (from left to right) under magnetic field directions of +B (red), –B (blue), and 0B (black), respectively. At least 430 current-voltage curves were obtained for all samples, with approximately 10 measurements performed on each Au NC. The error bar was defined as the standard deviation of all current values measured at each magnetic field condition. The I-V characterizations are performed by cAFM.
@@
Having established the chiral origin of the observed MR, we proceeded to investigate the robustness of this chirality-driven effect under various conditions, including different NC sizes, magnetic field strengths and direction, as well as AFM forces, to rule out potential artifacts. Using L-Au NCs as a model system, we observed that the current through chiral Au NCs increases gradually as the NC size increases from 150.2 to 350.5 nm, consistent with previously reported transport behaviors of metal nanoparticles [36] (Extended Data Figures 12a and 13). Concurrently, the overall MR value decreases with increasing NC size, while the MR values under –B consistently remain significantly larger than those under +B (Extended Data Figure 12d), confirming the persistence of asymmetric MR across different NC sizes. In the magnitude of the asymmetric MR also varies with the magnetic field strength. For L-Au NCs, as the external magnetic field increases from 0 to 1000 Oe, the current at +1.0 V remains nearly constant for +B, while it first decreases and later saturates beyond 300 Oe for –B. Correspondingly, the MR under +B keeps constant while the MR under –B substantially increases from 0% to 4800% from 0 to 300 Oe (Inset in Figure 2e and Extended Data Figures 12b, 12e, and 14). It is noteworthy that when B is perpendicular to the current direction, resistance of chiral Au NCs remains nearly identical to that of 0B (Extended data Figure 15). In cAFM measurements, the contact between the AFM tip and Au NCs sensitively influences the total resistance [37, 38]. To exclude potential artifacts arising from this contact, we evaluated the MR of L-Au NCs under various AFM forces. As anticipated, the current values increase with incremental force due to the compression of the molecular layer and the increased contact area between the tip and the Au NC. Notably, the asymmetric MR remains robust across different applied forces, although the MR at –B exhibits a weakly non-monotonic change with increasing force (Extended Data Figures 12c and 12f). This robustness against varying AFM force can be attributed to the stable morphology of crystalline nanoparticles compared to soft organic materials. In addition, we exclude that MR is caused by the displacement of Au NCs when applying the magnetic field. Because the magnetic susceptibility shows no chirality dependence (Extended Data Figure 5), the field-induced motion cannot generate chirality-dependent resistance even if it exists. As shown in the following Sec. 5, we further fabricated solid-state devices that fix all NCs and reproduced results measured by cAFM. These comprehensive investigations across different experimental parameters underscore the intrinsic nature and reliability of the observed chirality-driven magnetoresistance effect in Au NCs.
@@ 4 Control the sign of magnetoresistance
Figure 4: Magnetoresistance characterization of the sulfide-modified chiral Au NCs. (a) and (b) The experimental (dots) and fitted (line) current-voltage curves of Sm-L-Au and Sm-D-Au NCs at the magnetic field directions of +B (red), –B (blue), and 0B (black). We measured 180–650 current-voltage curves from more than 20 individual Au NCs and estimated the error bar by the standard deviation. (c) and (d) The MR of Sm-L-Au and Sm-D-Au NCs changing with bias under different external magnetic field directions. The Sm-L-Au and Sm-D-Au NCs were prepared by treating the L-Au and D-Au NCs with sodium sulfide, respectively. The I-V characterizations are performed by cAFM.
@@
The MR is positive for L/D-Au NCs as shown in Figure 2. We discovered that the sign of MR is actually sensitive to the chemical composition of the surface molecular layer on chiral Au NCs. We further treated the chiral Au NCs with sodium sulfide, resulting in additional sulfide adsorption into the CTAB layers on the Au NCs [39]. For clarity, we designate the sulfide-modified L- and D-Au NCs as Sm-L-Au and Sm-D-Au NCs, respectively. X-ray photoelectron spectroscopy studies confirm the presence of sulfide on their surfaces (Extended Data Figure 16). SEM and spectra studies demonstrate that Sm-L-Au and Sm-D-Au NCs maintain similar chiral morphologies and optically chiral properties to their non-modified counterparts (Extended Data Figures 17 and 18). Dynamic light scattering reveals slight size decreases for Sm-L-Au and Sm-D-Au NCs, attributed to the partial desorption of CTAB upon sulfide modification (Extended Data Figure 19 and Extended Data Table 3).
@@
We investigated the MR of Sm-L-Au and Sm-D-Au NCs using the same cAFM setup (Figure 4 and Extended Data Figure 20). At zero magnetic field, sulfide-modified NCs exhibit significantly higher conductance than pristine NCs, likely due to the thinner CTAB molecular layer. Despite this conductance change, the asymmetric nature of MR persists. For Sm-L(D)-Au NCs, –B (+B) field dramatically reduces the resistance by several times, while +B (–B) barely affects it. In contrast to L(D)-Au NCs, Sm-L(D) NCs display negative MR, as shown in Figure 4. This sign change in MR suggests that the magnetotransport properties are intimately related to the characteristics of the molecular shell layer.
Figure 5: Magneto-transport and asymmetric MR for chiral Au NCs in solid-state device. (a) Left: schematic illustration of the Au/chiral Au NC monolayer/Au solid-state device. Right: cross-sectional image of the Au/chiral Au NC monolayer/Au junction. (b) and (c) The MR of L-Au and D-Au NC monolayers measured under magnetic field ranging between -1.0 to 1.0 T and temperature between 10 and 400 K. (d) The trends of MR with temperature for L-Au NC monolayer at -1.0 T (black) and D-Au NC monolayer at +1.0 T (red). (e) The trends of resistance with temperature for L-Au NC monolayer (black) and D-Au NC monolayer (red) at 0 T. Both the MR and resistance are recorded at -1.0 V.
5 Magnetoresistance in solid-state devices
For chiral molecules, cAFM measurements typically show high MR, whereas the corresponding devices exhibit significantly lower MR ($\sim 1$%) [40, 41, 42, 43]. In contrast, we can reproduce the large MR in solid-state devices similar to the cAFM results for chiral Au NCs. These devices enable us to perform systematic measurements at varying temperature and magnetic field.
The solid-state devices were constructed by preparing the Au NC assembled monolayers on a template-stripped Au electrode and placing a PDMS (polydimethylsiloxane)-embedded Au electrode on top of the Au NC monolayers (Figure 5a and Extended Data Figure 21). The I-V curves through the Au NC monolayer junctions could be recorded under external magnetic fields and given temperature (Extended Data Figures 22 and 23). Similar to the results obtained by cAFM, the MR of chiral Au NC monolayers in solid-state devices exhibits a unidirectional response to the magnetic field direction. This MR response can be reversed by altering the chirality of Au NC and is independent of the applied order of the magnetic field (Figures 5b, 5c, and Extended Data Figures  24 and  25). The observed MR increases with magnetic field strength and saturates near 0.6 T for both L- and D-Au NC monolayers. The differences in MR saturation points between the solid-state device and cAFM measurements can be attributed to the number of chiral Au NCs involved and the types of electrodes used in these two methods. The cAFM measurement was conducted at single-particle level. In contrast, the contact area of the device is approximately 20,000 µm2, encompassing nearly one million Au NCs. A larger contact area introduces more defects, such as pin-hole or aggregation, causing the MR to reach saturation at higher magnetic fields. The different types of electrode contacts lead to various electrode-Au NC interfaces. As a control, the current of A-Au NC monolayer remains unchanged with the magnetic field direction even in the presence of a strong magnetic field (Extended Data Figures  24 and  26). More importantly, The current through both L- and D-Au NC monolayer junctions are temperature-independent, regardless of whether a magnetic field is applied. Correspondingly, the MR of L- and D-Au NC monolayers remains unaffected by temperature (Figures 5b-5e). These observation indicate a tunneling-dominated charge transport mechanism. Additional AC lock-in measurements of MR for both L- and D-Au NC monolayers corroborate the above results (Extended Data Figure 27).